Reduce the following fractions to the simplest form: 
 150 / 60 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

150 / 60 = (30× 5) / (30 × 2)
= 5/2

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The following fraction 150/60 can be reduced as
 
150/60
= (150÷30) / (60÷30)
= 5/2
 
Hope it helps!
Thank you